# Artificial Precision Polarization Array: Sensitivity for the axion-like dark matter with clock satellites

### v1

@@ Abstract
The approaches to searching for axion-like signals based on pulsars include observations with pulsar timing arrays (PTAs) and pulsar polarization arrays (PPAs). However, these methods are limited by observational uncertainties arising from multiple unknown and periodic physical effects, which substantially complicate subsequent data analysis. To mitigate these issues and improve data fidelity, we propose the Artificial Pulsar Polarization Arrays (APPA): a satellite network comprising multiple pulsed signal transmitters and a dedicated receiver satellite. In order to constrain the axion-photon coupling parameter $g_{a\gamma}$, we generate simulated observations using Monte Carlo methods to investigate APPA’s sensitivity via two complementary approaches: Bayesian analysis and frequentist analysis. Simulations indicate that for axion mass $m_{a}\sim\mathcal{O}\big(10^{-22}-10^{-19}\big)$ eV, APPA yields a better upper limit on $g_{a\gamma}$ (at the 95% confidence level) than conventional ground-based observations and achieves better detection sensitivity.
@@ I Introduction
The strong CP problem in quantum chromodynamics was famously addressed by Peccei and Quinn, who postulated a new $U_{PQ}(1)$ symmetry [10.1103/PhysRevLett.38.1440, 10.1103/PhysRevD.16.1791]. Now we know that the $U_{PQ}(1)$ breaking yields a pseudo Nambu–Goldstone boson: the axion [10.1103/PhysRevLett.40.223, 10.1103/PhysRevLett.40.279, 10.1088/1126-6708/2006/06/051]. Weinberg and Wilczek identified this axion as a light particle whose couplings could make it a viable dark matter (DM) candidate [10.1016/0370-2693(83)90637-8, 10.1016/0370-2693(83)90638-X]. In fact, ultralight axion-like particles (ALPs) arise generically in many high energy models and act as cold bosonic fields on cosmic scales [2203.14923, 10.1146/annurev-nucl-120720-031147, 10.1103/PhysRevD.95.043541]. These ultralight ALPs (with masses $m_{a}\ll 10^{-20}$eV) share the property that their de Broglie wavelength in galaxies is astrophysically large, implying a large occupation number in galactic halos and a collective wave like behavior [2507.00705, 10.1146/annurev-astro-120920-010024]. It was recognized that such fuzzy DM [10.1103/physrevlett.85.1158] could address apparent discrepancies of the collisionless cold DM paradigm on small scales. In particular, it was proposed that a boson of mass $m_{a}\sim\mathcal{O}\big(10^{-22}\big)$eV would stabilize gravitational collapse below kpc scales, yielding diffuse halo cores and strongly suppressing subgalactic power [10.1103/physrevlett.85.1158]. These wave effects, originating from the effective “quantum pressure” of the scalar field, modify the linear matter power spectrum and produce solitonic cores at halo centers. Subsequent work has developed the fuzzy DM model into a rich phenomenology [2507.00705].
The cosmological implications of ultralight dark matter (ULDM) are now tightly constrained by observations [10.1103/PhysRevD.95.043541, 10.1007/s00159-021-00135-6, 2203.14915]. In particular, Planck/WMAP data, combined with other CMB data constrain that axions with $m_{a}\lesssim 10^{-24}$eV contribute at most a few percent of the DM density (at 95% C.L.) [10.1103/PhysRevD.91.103512]. Likewise, analyses of the Lyman-$\alpha$ forest, which trace the small scale matter power spectrum, have dramatically improved the sensitivity [10.1103/PhysRevLett.119.031302, 10.1016/j.physrep.2016.06.005]. For example, it was reported that ULDM with the canonical $m_{a}\sim\mathcal{O}\big(10^{-22}\big)$eV is essentially excluded: their Lyman-$\alpha$ data imply $m_{a}>2\times 10^{-20}$eV at 95% C.L., thereby ruling out the mass range motivated by the string “axiverse” and small scale issues [10.1103/physrevlett.126.071302]. These cosmological bounds now push viable ULDM models to higher masses or subdominant fractions of the total DM [10.1103/PhysRevD.96.123514]. Complementary large scale surveys are beginning to probe intermediate mass ranges: recent analysises of eROSITA galaxy cluster counts and weak lensing data have set upper bounds on the axion relic density as low as $\Omega_{a}\lesssim 10^{-3}$ for $m_{a}\sim\mathcal{O}\big(10^{-27}–10^{-26}\big)$eV [2502.03353], tightening constraints on the relic fraction across the ultralight regime.
Astrophysical probes on galactic scales also impose stringent limits. Many attempts have been made to use stellar kinematics in dwarf spheroidal galaxies and low surface brightness halos to test the soliton core predictions of fuzzy DM [10.1038/nphys2996, 10.1103/PhysRevLett.113.261302, 10.1093/mnras/stx1887, 10.1103/physrevlett.123.051103, 10.3847/2041-8213/abf501, 10.1051/0004-6361/202346686]. For instance, a careful Jeans analysis of Fornax and Sculptor finds that if their halos were undisturbed by baryons, the core radii must exceed approximately 1 kpc and the axion mass would be constrained to $m_{a}<4\times 10^{-23}$eV (97.5% C.L.) in order to fit the data [10.1093/MNRAS/STX1941]. Other studies of rotation curves and stellar streams likewise seek evidence of the characteristic interference patterns or central cores induced by ULDM, but so far no convincing detection has been made [10.1103/physrevd.105.083015, 10.1051/0004-6361/202346686, 1808.00464, 2405.19410, 10.1088/1475-7516/2025/06/050].
On a different front, observations of compact objects provide powerful constraints on light bosons via superradiance. Rapidly spinning black holes, ranging from those in X-ray binaries to those detected in LIGO/Virgo events, would lose their spin through the superradiant growth of an axion cloud. Thus, the mere existence of high spin black holes rules out ALPs in specific mass windows [10.1007/978-3-030-46622-0, 10.1103/physrevd.83.044026, 10.1103/physrevd.96.035019, 2507.21788, 10.1103/PhysRevD.107.063021]. Recent analyses of gravitational wave binary BH spin measurements (e.g., GW231123) extend superradiance limits down to axion masses $m_{a}\sim\mathcal{O}\big(10^{-13}–10^{-12}\big)$eV. More generally, the “Regge plane” of black hole spins versus mass now excludes large swaths of the ALP parameter space [2507.21788].
@@
Novel timing and polarization techniques provide additional search strategies. In [10.1088/1475-7516/2014/02/019], it was noted that an oscillating ULDM field in the Galaxy would induce a characteristic, nearly monochromatic perturbation in pulsar timing residuals. PPTA [10.1103/PhysRevD.98.102002], NANOGrav [10.1088/1475-7516/2020/09/036], and EPTA [10.1103/PhysRevD.111.062005] pulsar timing data have been used to look for this signature: the data lead to upper limits on the axion coupling and local density in the mass range $m_{a}\sim\mathcal{O}\big(10^{-23}–10^{-21}\big)$eV. Independently, ultralight axions coupled to photons can cause cosmic birefringence: coherent ALP oscillations rotate the plane of linear polarization of passing light. For example, the high polarization pulses emitted by millisecond pulsars have been exploited to constrain ALPs as a birefringent medium [10.1103/PhysRevD.111.062005], and recent analyses of polarized fast radio burst (FRB) data have set competitive limits on the ALP-photon coupling in the $m_{a}\sim\mathcal{O}\big(10^{-21}–10^{-20}\big)$eV range [10.1038/s42005-025-02045-w]. In parallel, the mechanism of arraying many polarized sources has emerged: in [10.1103/PhysRevD.111.062005], the idea of a “pulsar polarization array(PPA)” was introduced to statistically enhance sensitivity to axion birefringence.
@@
In this work, we explore a novel avenue to probe this frontier using next generation timing and polarization networks. In particular,the concept of Artificial Precision Timing Array (APTA) was recently proposed in [2401.13668], where a constellation of clock satellites acts as artificial pulsars, to detect decihertz gravitational waves. The APTA detects subtle variations in signal parameters that may be induced by decihertz gravitational waves through precise and controllable signals. This enables the identification of the presence of decihertz gravitational waves. Compared with naturally occurring astrophysical PTAs, the APTA is far less affected by environmental and astrophysical uncertainties, providing a significant advantage. By analogy, one can envision the Artificial Precision Polarization Array (APPA) of precisely calibrated sources to detect axion induced birefringence. The APPA can detects the possible existence of axion induced birefringence by transmitting precisely controlled pulse signals with well defined polarization parameters. In this work, we investigate how the APPA can be leveraged to constrain or discover ultralight axion DM, focusing on the unique parameter ranges and systematics accessible to such terrestrial–space hybrid experiments.
@@
This paper is organized as follows. In Sec. II, we provide a brief overview of the axion induced birefringence effect and the principle of detecting axion signals through the rotation of the linear polarization angle. In Sec. III, we introduce the APPA concept in detail, Bayesian analysis as described in Sec. III.1 and the Generalized Lomb–Scargle Periodogram (GLSP) as presented in Sec. III.2. These methods are employed to derive the 95% C.L. upper limit $g_{95\%}$ and to compute sensitivity curves for the axion–photon coupling parameter $g_{a\gamma}$. The conclusions are summarized in Sec. IV. Throughout this work, we adopt natural units with $c=\hbar=1$ and choose the metric signature as (,+,+,+)(-,+,+,+).
@@ II Axion induced birefringence effect
The scalar axion field $a$ with mass $m_{a}$ can couple to the electromagnetic field $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$, and its action can be written as $S=\int d^{4}x\sqrt{-g}\mathcal{L}$, where $\mathcal{L}$ denotes the Lagrangian density:
@@
Here $\widetilde{F}^{\mu\nu}\equiv\frac{1}{2}\varepsilon^{\mu\nu\rho\sigma}F_{\rho\sigma}$ represents the dual electromagnetic tensor, and $g_{a\gamma}$ is the axion coupling coefficient.
@@
$\partial_{\mu}F^{\mu\nu}=g_{a\gamma}\widetilde{F}^{\mu\nu}\partial_{\mu}a.$ (II.2)
The birefringence effect for the electromagnetic field induced by axion field can be expressed as [10.1103/PhysRevD.111.062005, 10.1088/1475-7516/2022/06/014],
@@
Here $\hat{k}$ denotes the unit wave vector, and the details of the derivation are given in Appendix A. In an axion background, electromagnetic propagation is dispersive: the left and right circularly polarized modes acquire a tiny frequency splitting. Consequently, axion induced birefringence produces a relative phase shift between the two chiral polarization components [10.1103/PhysRevD.111.062005, 10.1088/1475-7516/2022/06/014, 2412.02229],
$\Delta\phi=\frac{g_{a\gamma}}{2}\int_{t_{s}}^{t_{o}}\frac{da}{dt}dt=\frac{g_{a\gamma}}{2}\left(a_{o}-a_{s}\right),$ (II.4)
where $a_{s}$ and $a_{o}$ denote the axion field at the signal emission point and the signal reception point, respectively. Unlike Faraday rotation, which depends on the observing frequency and on the line of sight magnetic field encountered along the propagation path, the polarization plane rotation induced by axion birefringence is determined by the axion field values at the endpoints. Consequently, the axion induced rotation is monochromatic.
In [10.1088/1475-7516/2014/02/019], the authors considered the occupation number of a scalar axion field and estimated it under standard local halo assumptions. Assuming a characteristic dark-matter velocity of $10^{-3}c$, and a local dark matter energy density $\rho_{\text{DM}}=0.4$GeV/cm3 in the vicinity of the Solar System, the occupation number can be estimated by
@@
$a\big(\vec{x},t\big)=a_{0}\big(\vec{x}\big)\mathrm{cos}\big(m_{a}t+\varphi\big(\vec{x}\big)\big),$ (II.6)
the amplitude of the oscillating scalar axion field is determined by the local dark matter energy density [10.1088/1475-7516/2019/02/059]:
@@
Meanwhile, noting that the typical dark matter velocity is $v\sim 10^{-3}c$ in realistic situations, the coherence time of the axion field can be expressed as $\tau_{c}=\frac{1}{m_{a}v^{2}}$. If the condition $\tau_{c}\gg T_{\text{obs}}$ is satisfied for a given axion mass, then the observed axion field amplitude behaves stochastically. As shown in [10.1103/physrevd.97.123006, 10.1038/s41467-021-27632-7], within a single coherence domain the amplitude distribution of the axion field follows a Rayleigh distribution,
$a_{0}=\frac{\sqrt{2\rho_{\text{DM}}}}{m_{a}}\alpha,\quad p\big(\alpha\big)\sim\alpha e^{-\frac{\alpha^{2}}{2}}.$ (II.8)
@@
$\phi_{a}$ $=\frac{g_{a\gamma}}{\sqrt{2}m_{a}}\left(\rho_{o}\alpha_{o}^{2}+\rho_{s}\alpha_{s}^{2}-2\sqrt{\rho_{o}\rho_{s}}\alpha_{o}\alpha_{s}\mathrm{cos}\Delta\right)^{\frac{1}{2}},$ (II.10)
with $\Delta=m_{a}T+\varphi_{o}-\varphi_{s}$, $\rho_{s}$ and $\rho_{o}$ denote the dark matter energy density in the vicinity of the source and the Solar System, respectively.
$T\equiv L/c$ is the signal travel time from a transmitting satellite to the receiving satellite.
@@ III Methods of analysis
1. Wide distance distribution of pulsars. The known pulsar population is predominantly concentrated within distances of a few kpc [10.1051/0004-6361/202243305]. The nearest pulsars lie at 120–170 pc, such as PSR J0437–4715 at about 157 pc [10.1086/529576] and PSR J2144–3933 at about 170 pc [10.1111/j.1745-3933.2011.01009.x]. Over such large propagation distances, pulsar signals are affected by a variety of complex astrophysical processes that tend to obscure potential axion induced signatures. These effects collectively complicate the modeling and degrade the achievable signal-to-noise ratio (S/N).
@@
3. Randomness of pulsar signal phase. The pulse phase $\varphi_{s}$ of a pulsar signal is inherently random and indeterminate, which prevents the signal propagation time $T$ from being used to reliably determine the value of $\Delta$ in Eq. (II.10).
4. Complex and frequency-dependent ionospheric interference. In practical observations, most pulsar signals are detected in the radio band, and ground-based radio telescopes therefore serve as the primary observational instruments. However, such measurements are inevitably affected by ionospheric interference. The Earth’s ionosphere induces Faraday rotation, which modifies all pulsar signals and leaves a characteristic imprint in the data. Consequently, this effect must be corrected during data processing. Although physical models of the Earth’s ionosphere have been extensively developed [10.1051/0004-6361/201220728, 10.1093/mnras/sty3324, 10.1029/2012RS004992], and a variety of algorithms for mitigating ionospheric contamination have been proposed [10.1086/501444, 10.1051/0004-6361/201220728, 10.1007/s00190-008-0266-1], these corrections nonetheless impose an additional computational burden.
To address these observational challenges at their source, we extend the concept proposed in [2401.13668] by introducing the APPA, a network of space-based satellites equipped with ultra precise clocks and capable of regularly transmitting pulsed signals to a dedicated receiving satellite. The transmitted signals are strictly periodic, with a timing uncertainty of $\sigma=\xi_{\delta t}\delta t$ [2401.13668], and carry preconfigured polarization information. As illustrated in Fig. 1, to ensure orbital stability, reduce operational costs, and minimize interference from the Sun and planets, the satellite network is designed to operate near the Sun–Earth Lagrange point $L_{2}$. The receiving satellite, positioned at the center of the transmitting network, continuously collects APPA signals and can accurately measure both the distance to and the direction of each transmitting satellite.
@@
$\lambda^{2}=\big(EF\sigma_{\text{err}}\big)^{2}+EQ^{2},$ (III.1)
here $\sigma_{\text{err}}$ denotes the instrumental measurement error, which we take to be $\sigma_{\text{err}}=0.3^{\circ}$. The $EF$ and $EQ$ represent the white noise scaling factor and the additive white noise, respectively, and their prior distributions are specified as follows:
@@
III.1 Bayesian analysis
In this section, we employ the parameters of the satellite network defined under the APPA concept and apply a Bayesian probabilistic framework to compute the 95% C.L. upper limit of $g_{a\gamma}$.
According to [10.1103/physrevlett.130.121401], the likelihood function that incorporates the axion signal is given by
$\mathcal{L}\big(\mathbf{\Theta}|\mathbf{d}\big)=\frac{1}{\sqrt{\mathrm{det}\big(2\pi\mathbf{\Sigma}\big)}}\mathrm{exp}\big(-\frac{1}{2}\mathbf{d^{T}}\cdot\mathbf{\Sigma}^{-1}\cdot\mathbf{d}\big),$ (III.3)
where $\mathbf{d}=\mathbf{s}+\mathbf{n}$ consists of the axion signal $\mathbf{s}$ and noise $\mathbf{n}$, $\mathbf{\Theta}=\{g,m_{a}\}$ denotes the set of model parameters, and $\mathbf{\Sigma}=\mathbf{\Sigma}^{(s)}+\mathbf{\Sigma}^{(n)}$ is the cross-correlation matrix. In practice, each satellite in the network typically produces on the order of $10^{4}\sim 10^{5}$ signals. As a result, $\mathbf{\Sigma}$ in Eq. (III.3) becomes prohibitively large, and the associated computational burden severely degrades the efficiency of solving the system. Fortunately, [10.1140/epjc/s10052-011-1554-0, 10.1103/PhysRevD.103.076018] demonstrate that $g_{95\%}$ can be determined from the condition: $\big\langle q\big\rangle=2.71$, where

### v2

@@ Abstract
The approaches to searching for axion-like signals based on pulsars include observations with pulsar timing arrays (PTAs) and pulsar polarization arrays (PPAs). However, these methods are limited by observational uncertainties arising from multiple unknown and periodic physical effects, which substantially complicate subsequent data analysis. To mitigate these issues and improve data fidelity, we propose the Artificial Pulsar Polarization Arrays (APPA): a satellite network comprising multiple pulsed signal transmitters and a dedicated receiver satellite. To constrain the axion-photon coupling parameter $g_{a\gamma}$, we generate simulated observations using Monte Carlo methods and investigate the sensitivity of APPA using two complementary approaches: Likelihood analysis and frequentist analysis. Simulations indicate that for the axion mass range of $10^{-22}-10^{-18}$ eV, APPA yields a tighter upper limit on $g_{a\gamma}$ (at the 95% confidence level) than conventional ground-based observations, while also achieving superior detection sensitivity. Moreover, a larger spatial distribution scale of the satellite network corresponds to a greater advantage in detecting axions with lighter masses.
@@ I Introduction
The strong CP problem in quantum chromodynamics was famously addressed by Peccei and Quinn, who postulated a new $U_{PQ}(1)$ symmetry [41, 40]. Now we know that the $U_{PQ}(1)$ breaking yields a pseudo Nambu–Goldstone boson: the axion [55, 56, 50, 9]. The axion can also be identified as a light particle whose couplings could make it a viable dark matter (DM) candidate [45, 15, 1]. The ultralight ALPs (with masses $m_{a}\ll 10^{-20}$eV) share the property that their de Broglie wavelength in galaxies is astrophysically large, implying a large occupation number in galactic halos and a collective wave like behavior [17, 28, 32]. It was recognized that such fuzzy DM [26] could address apparent discrepancies of the collisionless cold DM paradigm on small scales. In particular, it was proposed that a boson of mass $m_{a}\sim\mathcal{O}\big(10^{-22}\big)$eV would stabilize gravitational collapse below kpc scales, yielding diffuse halo cores and strongly suppressing subgalactic power [26]. These wave effects, originating from the effective “quantum pressure” of the scalar field, modify the linear matter power spectrum and produce solitonic cores at halo centers.
Cosmological observations provide important and effective methods of constraining ultralight dark matter (ULDM) [27, 18, 6]. Analyses of the Planck/WMAP data indicate that axions with $m_{a}\lesssim 10^{-24}$eV may constitute a fraction of the dark matter density (at 95% C.L.) [25]. Tracking the small scale matter power spectrum with the Lyman-$\alpha$ forest has significantly improved the sensitivity of axion searches [29, 35] and has excluded portions of the allowed axion mass range [46]. On galactic scales, astrophysical observations also place stringent limits on axions. For example, stellar kinematic analyses of dwarf spheroidal galaxies and low surface brightness halos can test the solitonic cores predicted by fuzzy dark matter [47, 48, 36, 34, 24, 7]; jeans analyses of observational data from different galaxies further constrain the mass range of axions and provide estimates of the core radii of dark matter halos [22]. Additional approaches for probing ULDM include searches for ULDM induced central cores using rotation curves and stellar streams [8, 7, 4, 11, 10], as well as constraints on the upper limit of the axion density derived from eROSITA galaxy cluster counts and weak lensing data [58].
@@
Novel timing and polarization techniques provide additional search strategies. In [32], it was noted that an oscillating ULDM field would induce a characteristic, nearly monochromatic perturbation in pulsar timing residuals. PPTA [44], NANOGrav [31], and EPTA [43] pulsar timing data have been used to look for this signature: the data lead to upper limits on the axion coupling and local density in the mass range $m_{a}\sim\mathcal{O}\big(10^{-23}–10^{-21}\big)$eV. Independently, ultralight axions coupled to photons can cause cosmic birefringence: coherent ALP oscillations rotate the plane of linear polarization of passing light. Constraints on the axion–photon coupling from birefringence effects in polarized millisecond-pulsar pulses and polarized fast radio bursts (FRBs) are given in [43] and [54]. In parallel, the mechanism of arraying many polarized sources has emerged: in [33], the idea of a “pulsar polarization array(PPA)” was introduced to statistically enhance sensitivity to axion birefringence. For further advances on constraining axion parameters through polarization measurements of signals from black holes, please see [21].
@@
In this work, we explore a novel avenue to probe this frontier using next generation timing and polarization networks. In particular, the concept of Artificial Precision Timing Array (APTA) was recently proposed in [2], where a constellation of clock satellites acts as artificial pulsars, to detect decihertz gravitational waves. The APTA detects subtle variations in signal parameters that may be induced by decihertz gravitational waves through precise and controllable signals. This enables the identification of the presence of decihertz gravitational waves. Compared with naturally occurring astrophysical PTAs, the APTA is far less affected by environmental and astrophysical uncertainties, providing a significant advantage. By analogy, one can envision the Artificial Precision Polarization Array (APPA) of precisely calibrated sources to detect axion induced birefringence. The APPA can detects the possible existence of axion induced birefringence by transmitting precisely controlled pulse signals with well defined polarization parameters. In this work, we investigate how the APPA can be leveraged to constrain or discover ultralight axion DM, focusing on the unique parameter ranges and systematics accessible to such terrestrial–space hybrid experiments.
@@
This paper is organized as follows. In Sec. II, we provide a brief overview of the axion induced birefringence effect and the principle of detecting axion signals through the rotation of the linear polarization angle. In Sec. III, we introduce the APPA concept in detail, likelihood analysis as described in Sec. III.1 and the Generalized Lomb–Scargle Periodogram (GLSP) as presented in Sec. III.2. These methods are employed to derive the 95% C.L. upper limit $g_{95\%}$ and to compute sensitivity curves for the axion–photon coupling parameter $g_{a\gamma}$. The conclusions are summarized in Sec. IV. Throughout this work, we adopt natural units with $c=\hbar=1$ and choose the metric signature as (,+,+,+)(-,+,+,+).
@@ II Axion induced birefringence effect
The action can be written as $S=\int d^{4}x\sqrt{-g}\mathcal{L}$, where $\mathcal{L}$ denotes the Lagrangian density:
@@
The axion field $a$ with mass $m_{a}$ can couple to the electromagnetic field $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$, and $\widetilde{F}^{\mu\nu}\equiv\frac{1}{2}\varepsilon^{\mu\nu\rho\sigma}F_{\rho\sigma}$ represents the dual electromagnetic tensor, and $g_{a\gamma}$ is the axion coupling coefficient.
@@
$\partial_{\mu}F^{\mu\nu}=g_{a\gamma}\partial_{\mu}a\widetilde{F}^{\mu\nu}.$ (II.2)
The birefringence effect for the electromagnetic field induced by axion field can be expressed as [43, 12],
@@
Here $\hat{k}$ denotes the unit wave vector, and the details of the derivation are given in Appendix A. In an axion background, electromagnetic propagation is dispersive: the left and right circularly polarized modes acquire a tiny frequency splitting. Consequently, axion induced birefringence produces a relative phase shift between the two chiral polarization components [43, 12, 57],
$\Delta\phi=\frac{g_{a\gamma}}{2}\int_{t_{s}}^{t_{o}}\frac{da}{dt}dt=\frac{g_{a\gamma}}{2}\left(a_{o}-a_{s}\right).$ (II.4)
Here, $a_{s}$ and $a_{o}$ denote the axion field at the signal emission point and the signal reception point, respectively. Unlike Faraday rotation, which depends on the observing frequency and on the line of sight magnetic field encountered along the propagation path, the polarization plane rotation induced by axion birefringence is determined by the axion field values at the endpoints. Consequently, the axion induced rotation is monochromatic.
In [32], the authors considered the occupation number of a scalar axion field and estimated it under standard local halo assumptions. Assuming a characteristic dark-matter velocity of $10^{-3}c$, and a local dark matter energy density $\rho_{\text{DM}}=0.4$GeV/cm3 in the vicinity of the Solar System, the occupation number can be estimated by
@@
$a\big(\vec{x},t\big)=a_{0}\big(\vec{x}\big)\mathrm{cos}\big[m_{a}t+\varphi\big(\vec{x}\big)\big].$ (II.6)
The amplitude of the oscillating scalar axion field is determined by the local dark matter energy density [30]:
@@
Meanwhile, noting that the typical dark matter velocity is $v\sim 10^{-3}c$. In realistic situations, the coherence time of the axion field can be expressed as $\tau_{c}=\frac{1}{m_{a}v^{2}}$. If the condition $\tau_{c}\gg T_{\text{obs}}$ is satisfied for a given axion mass, then the observed axion field amplitude behaves stochastically. As shown in [20, 13], within a single coherence domain the amplitude distribution of the axion field follows a Rayleigh distribution,
$a_{0}=\frac{\sqrt{2\rho_{\text{DM}}}}{m_{a}}\alpha,\qquad p\big(\alpha\big)\sim\alpha e^{-\frac{\alpha^{2}}{2}}.$ (II.8)
@@
$\phi_{a}$ $=\frac{g_{a\gamma}}{\sqrt{2}m_{a}}\left(\rho_{o}\alpha_{o}^{2}+\rho_{s}\alpha_{s}^{2}-2\sqrt{\rho_{o}\rho_{s}}\alpha_{o}\alpha_{s}\mathrm{cos}\theta_{T}\right)^{\frac{1}{2}},$ (II.10)
with $\theta_{T}=m_{a}T_{i}+\varphi_{o}-\varphi_{s}$, $\rho_{s}$ and $\rho_{o}$ denote the dark matter energy density in the vicinity of the source and the Solar System, respectively. $T_{i}\equiv L_{i}/c$ is the signal travel time from a transmitting satellite $S_{i}$ to the receiving satellite $O$.
@@ III Methods of analysis
1. Wide distance distribution of pulsars. The known pulsar population is predominantly concentrated within distances of a few kpc [16]. The nearest pulsars lie at 120–170 pc, such as PSR J0437–4715 at about 157 pc [53] and PSR J2144–3933 at about 170 pc [51]. Over such large propagation distances, pulsar signals are affected by a variety of complex astrophysical processes that tend to obscure potential axion induced signatures. These effects collectively complicate the modeling and degrade the achievable signal-to-noise ratio (S/N).
@@
3. Randomness of pulsar signal phase. The pulse phase $\varphi_{s}$ of a pulsar signal is inherently random and indeterminate, which prevents the signal propagation time $T$ from being used to reliably determine the value of $\Theta$ in Eq. (II.10).
4. Complex and frequency-dependent ionospheric interference. In practical observations, most pulsar signals are detected in the radio band, and ground-based radio telescopes therefore serve as the primary observational instruments. However, such measurements are inevitably affected by ionospheric interference. The Earth’s ionosphere induces Faraday rotation, which modifies all pulsar signals and leaves a characteristic imprint in the data. Consequently, this effect must be corrected during data processing. Although physical models of the Earth’s ionosphere have been extensively developed [49, 42, 38], and a variety of algorithms for mitigating ionospheric contamination have been proposed [23, 49, 39], these corrections nonetheless impose an additional computational burden.
To address these observational challenges at their source, we extend the concept proposed in [2] by introducing the APPA, a network of space-based satellites equipped with ultra precise clocks and capable of regularly transmitting pulsed signals to a dedicated receiving satellite. The transmitted signals are strictly periodic, with a timing uncertainty of $\sigma_{t}=\xi_{\delta t}\delta t$ [2], and carry preconfigured polarization information. As illustrated in Fig. 1, the receiving satellite, positioned at the center of the transmitting network, continuously collects APPA signals and can accurately measure both the distance to and the direction of each transmitting satellite.
@@
$\lambda=\big(EF\sigma_{\text{err}}\big)^{2}+EQ^{2}.$ (III.1)
Here, $\sigma_{\text{err}}$ denotes the instrumental measurement error, which we take to be $\sigma_{\text{err}}=0.3^{\circ}$. The $EF$ and $EQ$ represent the white noise scaling factor and the additive white noise, respectively, and their prior distributions are specified as follows:
@@
First, we show the main conclusions of this work in advance. We present a comparison between the simulation results of this work and existing observational constraints in a single figure (see Fig. 2). Owing to the long observational baseline $T_{\text{obs}}$ and the highly stable signal emission cadence $\Delta t$, APPA is theoretically capable of probing an exceptionally broad range of axion induced oscillation frequencies. However, as shown in Fig. 2, although APPA theoretically possesses an extremely wide detection range, the domain where it truly demonstrates an advantage in axion signal detection is limited to $m_{a}\sim\mathcal{O}\big(10^{-22}–10^{-18}\big)$eV, both the 95% confidence upper limit $g_{95\%}$ derived from likelihood analysis and the detection sensitivity $g_{a\gamma}$ obtained via frequentist analysis highlight the significant advantages of the APPA framework over current ground-based observations.
Figure 2: Comparison between the APPA simulation results and existing observational constraints. Color coding: red, CAST [5]; magenta, PPA [57]; green, PPTA [12]; blue, likelihood analysis (L=5L=5AU), we shows that APPA has a detection advantage within $m_{a}\sim\mathcal{O}\big(10^{-22}–10^{-18}\big)$eV and exhibits comparable sensitivity to axions with different masses; purple, frequentist analysis, the detection sensitivity is set to $\phi_{a}=0.01^{\circ}$ in Eq. (III.9), consistent with current observational limits on linear polarization angle rotation, with minor fluctuations attributed to instrumental noise. The black dashed line marks the critical boundary, to the left of this boundary, the condition $\mathrm{cos}\big(m_{a}T\big)\simeq 1$ holds true, to the right it is no longer valid.
III.1 Likelihood analysis
In this section, we employ the parameters of the satellite network defined under the APPA concept and apply the likelihood analysis to compute the 95% C.L. upper limit of $g_{a\gamma}$.
According to [33], the likelihood function that incorporates the axion signal is given by
[…]
where the observational data $\mathbf{d}=\mathbf{s}+\mathbf{n}$ consists of the axion signal $\mathbf{s}$ and noise $\mathbf{n}$. $\mathbf{\widetilde{d}}$ denotes $\mathbf{d}$ with the intrinsic effect subtracted, $\mathbf{\Theta}=\{g,m_{a}\}$ denotes the set of model parameters and $\mathbf{\Sigma}=\mathbf{\Sigma}^{(s)}+\mathbf{\Sigma}^{(n)}$ is the cross-correlation matrix. In practice, each satellite in the network typically produces on the order of $10^{4}\sim 10^{5}$ signals. As a result, $\mathbf{\Sigma}$ in Eq. (III.3) becomes prohibitively large, and the associated computational burden severely degrades the efficiency of solving the system. Fortunately, [33, 14, 19] demonstrate that $g_{95\%}$ can be determined from the condition: $\big\langle q\big\rangle=2.71$, where